# Difference Decomposition Networks for Infrared Small Target Detection

### v1

@@ Abstract
One of main challenges faced by infrared small target detection (ISTD) is weak targets. To enhance target saliency, recent studies have widely adopted difference-based methods, which highlight the weak target by computing and weighting difference features. However, these methods rely on fixed and predetermined weights to combine these difference features in spatial and temporal domains. These approaches fail to adaptively capture the inter-correlations among various directions or different frames. Moreover, these inter-correlations contain crucial structural information about the target. Consequently, this failure limits the model’s performance. To address this limitation, we propose a Basis Decomposition Module (BDM) to extract the correlation information in different directions. For single-frame ISTD (SISTD), we introduce Spatial Difference Decomposition Module (SD2M) and Spatial Difference Decomposition Downsampling Module (SD3M), all based on the BDM structure, to extract the detailed information in the spatial dimensions. In addition, we use the BDM-based Temporal Difference Decomposition Module (TD2M) to extract the motion information for multi-frame ISTD (MISTD). By incorporating our proposed modules into a stacked U-Net architecture, we introduce two networks: the Spatial Difference Network (SD2Net) for SISTD and the Spatio-Temporal Difference Network (STD2Net) for MISTD. These networks are designed to achieve precise detection of infrared small targets without a large number of parameters. On the SISTD task, SD2Net achieves competitive performance against most established networks with 0.27M parameters. On the MISTD datasets, STD2Net achieves a mIoU of 87.68%, outperforming SD2Net, which achieves a mIoU of 64.97%. Our codes are available: https://github.com/greekinRoma/IRSTD_HC_Platform.
@@ I Introduction
Unlike visible light imaging, infrared sensing maintains reliable performance in adverse weather conditions, offering high detection reliability, concealment, and mobility. As one of the main applications of infrared sensing, Infrared Search and Track (IRST) have demonstrated the strong robustness under varying circumstances, and works as the fundament in several applications [Zhang2024IRPruneDet, Zhu2025PFF, iMoPKL2025, MICPL2025].
Infrared small target detection (ISTD) is the core of IRST systems, and has attracted growing research interests. Nevertheless, the dim targets and complex backgrounds render ISTD significantly more difficult than many other detection tasks [Zhu2024GSTUnet]. Therefore, ISTD is a fundamental and challenging computer vision task, and merits further investigation [dai2025SeRankDet, zhang2025daaf, zhang2025coupled, zhang2025dual].
To address these challenges in ISTD, numerous model-based approaches are proposed for single-frame ISTD (SISTD) [kong2022LogTFNN, cao2021infrared, dai2017reweighted, wang2019miss]. With the recent development of deep-learning techniques, data-driven methods have demonstrated superior performance to model-driven methods [Zhang2025IRmamba, li2025hstnet, li2025mmidnet, zhang2025exploringstatespacemodel]. Up to now, the methods based on deep learning have become the mainstream of ISTD.
In recent years, difference-based methods, like central difference convolutions (CDCs) and difference-frame methods, have gained the increasing attention [Wu2025Salm]. In these modules, the detailed information are highlighted using high-pass kernels in spatial or temporal dimensions[Zhang2021ECA]. However, both CDC-based approaches and other spatial-domain methods face a fundamental limitation in distinguishing small targets from structurally similar clutter without the leverage of motion information.
In order to solve above problems, motion information is introduced into the algorithms. Existing general video object detection (GVOD) methods [Deng2019Object, zhu2017deep, wu2019sequence, zhou2022transvod, wang2022ptseformer] are designed for large targets with apparent inter-frame motion. Nevertheless, the infrared small targets are usually dim and move slowly in infrared streams[Duan2024Triple, zhang2025fsatfusion, zhang2025wife], and the GVOD methods could not highlight the motion information of dot targets. Therefore, GVOD approaches are not suitable for ISTD. This motivates the exploration of new algorithms designed for infrared small targets. Currently, the common technique for obtaining motion information for infrared small targets relies on the difference between key frames and reference frames [Li2025dtum].
As illustrated in Figure 1, difference-based methods all involve two steps: first, they extract difference features by subtracting, and then combine these features using coefficients. These coefficients are learned during training but remain static during inference.
@@
Figure 1: The standard CDC and frame-difference operations, both employing predetermined static coefficients to combine difference results.
According to basis decomposition, an input can be represented as a combination of basis features and their dynamic coefficients. These input-specific coefficients are unique per basis and inherently capture correlation information between features. To address the limitation of fixed weights in difference-based methods, we introduce basis decomposition. This enables adaptive feature combination based on the learned correlations, thereby improving performance.
Based on basis decomposition, we construct a unified and lightweight framework for ISTD, named the Basis Decomposition Module (BDM). The framework can enhance task-relevant information, like difference information, in both spatial and temporal domains. These coefficients can be regarded as an inherent property of the input features. Therefore, their extraction does not require any additional parameters.
Building on BDM, we design a series of networks collectively referred to as Difference Decomposition Networks (DDNets). For SISTD, we utilize adaptations of BDM in the spatial domain, specifically the Spatial Difference Decomposition Module (SD2M) and the Spatial Difference Decomposition Downsampling Module (SD3M), to enhance high-frequency target features while suppressing low-frequency backgrounds. Based on two modules, we propose the Spatial Difference Decomposition Network (SD2Net). For MISTD, we further extend SD2Net to Spatiotemporal Difference Decomposition Network (STD2Net) by integrating temporal information through the Temporal Difference Decomposition Module (TD2M), which emphasizes the differences among frames and the features of the current frame in the temporal domain. The proposed networks not only outperform existing methods but also have a smaller model size. The contributions are summarized as below:
@@
    We introduce the BDM, a foundational component that enhances task-relevant features while suppressing task-irrelevant information through basis decomposition.
@@
    We develop SD2M and SD3M, derived from BDM, into a modified U-Net, proposing SD2Net for SISTD to strengthen high-frequency target features and suppress low-frequency clutter with few parameters.
@@
    For detecting moving targets against static backgrounds, we propose the TD2M. By incorporating TD2M, we further build STD2Net, which effectively highlights moving targets and suppresses the clutter.
@@ II-A Single-frame Infrared Small Target Detection
Existing SISTD methods can be categorized into two methodological paradigms: model-driven methods and data-driven methods.
Model-driven methods leverage domain-specific priors and bifurcate into three branches. Filter-based methods [Qin2019Facetkernel] enhance target features and suppress clutter with filters.Although these techniques perform well in simple scenes, they are highly sensitive to window size and easily fail when clutter exhibits similar texture or brightness. Human visual system-based methods [WeiYouLi2016Multiscale, Han2021WLCM] are based on visual characteristics of the visual system.However, these methods rely on handcrafted contrast metrics and struggle to generalize across diverse imaging conditions. Optimization-based methods [gao2013infrared, ZhangPeng2019Infrared] formulate ISTD as an optimization problem to detect the targets by separating them from the background.Despite their interpretability, these models usually require iterative solvers, leading to heavy computation and limited scalability Although these methods are computationally efficient and interpretable, they are usually sensitive to parameter selection and fail under complex cluttered scenes.
Data-driven methods have addressed these issues of SISTD by learning discriminative features from large datasets. Dai et al. have successively proposed the Asymmetric Contextual Modulation (ACM) [DaiWuZhouBarnard2021Asymmetric] and Attentional Local Contrast Network (ALCNet) [dai2021attentional] for ISTD. Dense Nested Attention Network (DNANet) [RenLiHanShu2021DNANet] supports adaptive interactions between feature layers to prevent the loss of targets. Furthermore, UIUNet [WuHongChanussot2023UIU-Net] employs a simple and effective “U-Net in U-Net” framework to enhance both global and local contrast information. Receptive-field and Direction-induced Attention Network (RDIAN) [SunBaiYangBai2023Receptive-Field] utilizes the characteristics of targets to solve the imbalance between targets and background. Infrared Small-target Detection U-Net (ISTDU-Net) [HouZhangTanXiZhengLi2022ISTDU-Net] introduces a fully connected layer in the skip connection to suppress the backgrounds with similar structures from the global receptive field. The Attention-guided Pyramid Context Network (AGPCNet) [ZhangCaoPuPeng2021AGPCNet] employs an attention-guided context block, providing the model with a perspective on both inner and global patches. Spatial-Channel Cross Transformer Network (SCTransNet) [Yuan2024SCtransNet] uses the Transformer to obtain the effective global information.
Nevertheless, these architectures tend to be computationally heavy, and the lack of explicit mechanisms for separating target-relevant and background-redundant features often leads to false alarms under complex thermal clutter. Therefore, how to design a lightweight, interpretable, and generalizable representation mechanism remains an open issue in single-frame ISTD.
@@ II-B Multi-frame Infrared Small Target Detection
Pre-extraction frameworks extract motion features before a SISTD network. Energy Accumulation (IFEA) [Du2022IFEA] and Spatio-temporal Differential Multiscale Attention Network (STDMANet) [Yan2023STDMANet] enhance the energy of targets by computing the difference between the current frame and reference frames before input into networks.
Result-based extraction frameworks extract temporal information based on the results of single-frame neural networks. Direction-code Temporal U-shape Module (DTUM) [Li2025dtum] designs a new algorithm to extract the motion features from the results of SISTD networks. These approaches emphasize explicit motion cues and are straightforward to implement, but their reliance on handcrafted frame-difference operators makes them sensitive to background dynamics and sensor noise, often resulting in unstable motion estimation. The Recurrent Feature Refinement (RFR) framework [ying2025rfr] effectively captures long-term temporal dependencies based on the outputs of SISTD networks. Although such methods exploit temporal refinement after initial detection, they often inherit errors from the single-frame stage, causing accumulated false alarms when the early predictions are unreliable.
@@
Spatiotemporal feature simultaneous extraction frameworks obtain the temporal and spatial information using modules in a single network. Sliced Spatio-temporal Network (SSTNet) [Chen2024SSTNet] utilizes the ConvLSTM node to extract the motion information of the small infrared targets. The Lightweight Asymmetric Spatial Feature Network (LASNet) [Chen2024LASNet] utilizes motion feature extraction and motion-affinity fusion for extracting temporal information. The Triple-domain Strategy (Tridos) [Duan2024Triple] employs frequency-aware enhancement to enhance the detection of dim infrared targets. The Local Motion Aware Transformer (LMAFormer) [huang2024lmaformer] utilizes local motion-aware attention to enhance the weak target motion information in MISTD. These end-to-end architectures can effectively integrate spatial and temporal cues, yet they usually demand heavy computation and high memory usage, limiting their deployment in real-time infrared tracking systems. Furthermore, their focus on global aggregation often overlooks feature disentanglement, making small targets difficult to separate from cluttered backgrounds.
Although these multi-frame networks improve robustness, most rely on heuristic motion aggregation or fusion-based temporal modules. They lack a unified and interpretable mechanism to decouple target-relevant and background-redundant components, leading to limited generalization across different ISTD tasks.
In summary, current SISTD and MISTD methods are largely designed for specific subtasks, lacking theoretical extensibility and efficiency. To overcome these limitations, we introduce a Basis Decomposition Module (BDM) that provides a unified, lightweight foundation for both spatial and temporal feature enhancement.
@@ III-A Preliminaries
((a))
((b))
Figure 2: (a) The procedure of basis decomposition. (b) The illustration of the proposed framework for signal recovery and task-relevant feature enhancement. First, we collect both task-relevant and task-irrelevant basis features into a basis pool according to the certain tasks. Then, the input is then processed in two parallel branches for recovery and enhancement. The upper branch indicates the recovery process, where the input is reconstructed through basis decomposition. The lower branch outlines the enhancement procedure, which selects the task-relevant basis features to reconstruct the input.
@@
As shown in the upper part of Figure 2, a vector $\vec{x}\in\mathbb{R}^{c_{\text{ful}}}$ can be recovered by $c_{ful}$ orthonormal vectors:
@@
where $s_{i}$ denotes the coefficient of $\vec{x}$ projected onto the basis vector $\vec{b}_{i}$. $\top$} is the transposition operation, and $c_{\text{ful}}$ is the dimension of the input vector.
However, by removing redundant basis vectors, the result is constructed solely from the selected basis elements, thereby enhancing the task-relevant features, as illustrated in the lower portion of Figure 2. By retaining only those vectors that are relevant to the task and removing irrelevant ones, we can strengthen the important features while reducing interference from non-essential information.
@@
In addition, as shown in Table I, our method does not require the basis vectors to be strictly orthonormal. This relaxation of the orthogonality constraint not only simplifies the computation process but also improve the performance.
When it comes to ISTD dependent on comprehensive and detailed information, we could select the difference information to construct the feature. Consequently, we retain the essential features as basis feature vectors to form $\vec{y}$ and approximate $\vec{x}$.

### v2

@@ Abstract
Weak target signal response is one of the key challenges in infrared small target detection (IRSTD), leading to missing detections, as some dim targets can be completely overwhelmed by the background. Recent studies have widely adopted difference-based methods, which utilize difference features in spatial and temporal domains. These methods rely on fixed and predetermined coefficients to combine these difference features and prevent the targets from overwhelmed by complex background, but fail to extract the correlation information between difference features and original information. To address this, we propose the Basis Decomposition Module (BDM) that can extract correlation information. Building on BDM, we develop the several variants for IRSTD and name them Difference Decomposition Modules (DDMs), which treat the normalized difference results as the basis features, combine them with the coefficient of these features projected onto the input, and achieve precise detection of small infrared targets without requiring a large number of parameters. Besides target enhancement from difference operations, DDMs also demonstrate robustness to diverse transformations. By integrating DDMs into U-shaped architectures, we develop two networks: the Spatial Difference Network (SD2Net) and the Spatiotemporal Difference Network (STD2Net) for single-frame IRSTD (SIRSTD) and multi-frame IRSTD (MIRSTD), respectively. With DDMs, they demonstrate performance competitive with most established networks on the SIRSTD and MIRSTD tasks. Our codes will be made public at: https://github.com/greekinRoma/IRSTD_HC_Platform.
@@ I Introduction
Unlike visible light imaging, infrared sensing maintains reliable performance in adverse weather conditions, offering high detection reliability, concealment, and mobility. As one of the primary applications of infrared sensing, Infrared Search and Track (IRST) has demonstrated strong robustness under varying circumstances and serves as the foundation for several applications [Zhang2024IRPruneDet, Zhu2025PFF, iMoPKL2025, MICPL2025, yuan2025ASCNet]. Infrared small target detection (IRSTD) is a crucial task in IRST systems, garnering growing research interest. Nevertheless, the dim targets and complex backgrounds render IRSTD significantly more difficult than many other detection tasks [Zhu2024GSTUnet]. Therefore, IRSTD is a fundamental and challenging computer vision task, and merits further investigation [dai2025SeRankDet, zhang2025daaf, zhang2025coupled, yuan2025EDGSP].
In the past years, numerous model-based approaches have been proposed for single-frame IRSTD (SIRSTD) [kong2022LogTFNN, cao2021infrared, dai2017reweighted, wang2019miss]. Thanks to advances in deep learning, data-driven methods have become dominant in IRSTD, outperforming model-driven approaches [Zhang2025IRmamba, li2025hstnet, li2025mmidnet, zhu2025shifting]. However, spatial-based methods perform poorly when targets exhibit substantial visual similarity to their surroundings. In such circumstances, the primary difference between targets and false alarms is their motion. Thus, it is essential to introduce the motion information in complex backgrounds.
@@
Figure 1: The standard CDC and frame-difference operations, both employing predetermined static coefficients to combine difference results. (a) We employ difference kernels to obtain difference features across various directions and combine them with static weights. (b) We extract motion features by subtracting the key frame from the reference frames and fuse them with static coefficients.
There are some general video object detection (GVOD) methods classifying targets from their false alarms [Deng2019Object, zhu2017deep, wu2019sequence, zhou2022transvod, wang2022ptseformer]. The infrared small targets are usually dim and move slowly in infrared streams [Duan2024Triple, zhang2025fsatfusion, zhang2025wife], but the GVOD methods could not highlight the motion information of such targets. Therefore, GVOD approaches are not suitable for IRSTD, motivating the exploration of new algorithms designed for infrared small targets. Currently, the standard technique for obtaining motion information for infrared small targets relies on the difference between key frames and reference frames [Li2025dtum].
In recent years, difference-based methods, such as central difference convolutions (CDCs) [Wu2025Salm] and difference-frame methods [Yan2023STDMANet], have gained an increasing attention. In these modules, detailed information is highlighted using high-pass kernels in either spatial or temporal dimensions [Zhang2021ECA]. However, both CDC-based approaches and other difference-based methods face a fundamental limitation in static coefficients, as illustrated in Figure 1.
From the perspective of basis decomposition, an input feature vector can be represented as a linear combination of a set of basis vectors, or basis features, with their corresponding coefficients. These coefficients are computed by projecting the basis features onto the input, thereby encoding the correlation between each basis feature and the input itself. We can exploit these coefficients to reconstruct the input, when the basis features are enough. If certain basis features are not enough, the input may lose the information those missed features represent,and highlight the retained ones.
Building on this idea, we design a unified and lightweight module for IRSTD, referred to as the Basis Decomposition Module (BDM). BDM selects specific features as basis features based on the given task, performs basis decomposition on the input using these features to obtain projection coefficients, and generates a feature as a combination of the weights and features. In this way, we could utilize different features as basis features for varied tasks.
For IRSTD, we modify the BDM and develop the Difference Decomposition Modules (DDMs) by utilizing the normalized difference results as our basis features. The output of each DDM is a combination of difference features, and thus strengthen the dim targets. Additionally, this design has two key properties that could enhance the robustness of our module:
  • (1)
    The normalization and difference process removes absolute intensity information, ensuring that the features only retain relative information. In this way, the module focuses on relative changes, unaffected by variations in the overall intensity of the input.
  • (2)
    The weights are dynamic and adjust based on the input. For example, when the input rotates in the spatial domain, the weights change accordingly. This design is particularly effective in space-based scenarios, such as satellite imagery, where perspective rotations are common. Therefore, this design makes it more flexible when dealing with spatial and temporal variations.
We introduce DDMs to a network and propose the Spatial Difference Network (SD2Net) and the Spatiotemporal Difference Network (STD2Net) for SIRSTD and MIRSTD, achieving a robust performance in their corresponding tasks using a limited number of parameters. To sum up, the contributions of this work are summarized as follows:
@@
    We introduce the BDM to enhance task-relevant features and suppress irrelevant information through basis decomposition, and thereby propose the DDMs for IRSTD.
@@
    We develop the SD2Net for SIRSTD to enhance high-frequency target features and suppress low-frequency clutter with a minimal number of parameters.
@@
    We further build the STD2Net, which is based on the SD2Net and effectively highlights moving targets while suppressing the static clutter for MIRSTD.
@@ II-A Single-frame Infrared Small Target Detection
Existing SIRSTD methods can be categorized into two methodological paradigms: model-driven methods and data-driven methods.
Model-driven methods leverage domain-specific priors and bifurcate into three branches. Filter-based methods [Qin2019Facetkernel] enhance target features and suppress clutter with filters. Although these techniques perform well in simple scenes, they are susceptible to variations in window size and can easily fail when clutter exhibits similar textures or brightness. Human visual system-based methods [WeiYouLi2016Multiscale, Han2021WLCM] are based on visual characteristics of the visual system. However, these methods rely on handcrafted contrast metrics and struggle to generalize across diverse imaging conditions. Optimization-based methods [gao2013infrared, ZhangPeng2019Infrared] formulate IRSTD as an optimization problem to detect the targets by separating them from the background. Despite their interpretability, these models typically require iterative solvers, resulting in significant computation and limited scalability. Although these methods are computationally efficient and interpretable, their performance is often sensitive to parameter selection and unstable in complex, cluttered scenes.
Data-driven methods have addressed these issues of SIRSTD by learning discriminative features from large datasets. Dai et al. have successively proposed the Asymmetric Contextual Modulation (ACM) [DaiWuZhouBarnard2021Asymmetric] and Attentional Local Contrast Network (ALCNet) [dai2021attentional] for IRSTD. Dense Nested Attention Network (DNANet) [RenLiHanShu2021DNANet] supports adaptive interactions between feature layers to prevent the loss of targets. Furthermore, UIUNet [WuHongChanussot2023UIU-Net] employs a simple and effective “U-Net in U-Net” framework to enhance both global and local contrast information. Receptive-field and Direction-induced Attention Network (RDIAN) [SunBaiYangBai2023Receptive-Field] utilizes the characteristics of targets to solve the imbalance between targets and background. Infrared Small-target Detection U-Net (ISTDU-Net) [HouZhangTanXiZhengLi2022ISTDU-Net] introduces a fully connected layer in the skip connection to suppress the backgrounds with similar structures from the global receptive field. The Attention-guided Pyramid Context Network (AGPCNet) [ZhangCaoPuPeng2021AGPCNet] employs an attention-guided context block, providing the model with a perspective on both inner and global patches. The Spatial-Channel Cross Transformer Network (SCTransNet) [Yuan2024SCtransNet] utilizes the Transformer to extract practical global information.
Whereas previous models focused primarily on precision, our approach places greater emphasis on lightweight network design, aiming to strike a balance between computational cost and performance. This enables the single-frame model to remain sufficiently lightweight, creating favorable conditions for incorporating temporal information.
@@ II-B Multi-frame Infrared Small Target Detection
Pre-extraction frameworks extract motion features before a SIRSTD network. Energy Accumulation (IFEA) [Du2022IFEA] and Spatiotemporal Differential Multiscale Attention Network (STDMANet) [Yan2023STDMANet] enhance the energy of targets by computing the difference between the current frame and reference frames before input into networks.
@@
Spatiotemporal feature simultaneous extraction frameworks obtain the temporal and spatial information using modules in a single network. Sliced Spatiotemporal Network (SSTNet) [Chen2024SSTNet] utilizes the ConvLSTM node to extract the motion information of the small infrared targets. The Lightweight Asymmetric Spatial Feature Network (LASNet) [Chen2024LASNet] utilizes motion feature extraction and motion-affinity fusion for extracting temporal information. The Triple-domain Strategy (Tridos) [Duan2024Triple] employs frequency-aware enhancement to enhance the detection of dim infrared targets. The Local Motion Aware Transformer (LMAFormer) [huang2024lmaformer] utilizes local motion-aware attention to enhance the weak target motion information in MIRSTD. These end-to-end architectures can effectively integrate spatial and temporal cues; however, they typically require heavy computation and high memory usage, which limits their deployment in real-time infrared tracking systems. Furthermore, their focus on global aggregation often overlooks feature disentanglement, making it challenging to separate small targets from cluttered backgrounds.
Result-based extraction frameworks extract temporal information based on the results of single-frame neural networks. Direction-code Temporal U-shape Module (DTUM) [Li2025dtum] designs a new algorithm to extract the motion features from the results of SIRSTD networks. These approaches emphasize explicit motion cues and are straightforward to implement. Still, their reliance on handcrafted frame-difference operators makes them sensitive to background dynamics and sensor noise, often resulting in unstable motion estimation. The Recurrent Feature Refinement (RFR) framework [ying2025rfr] effectively captures long-term temporal dependencies based on the outputs of SIRSTD networks.
We adopt the result-based extraction framework as our foundation because its decoupling from feature-level fusion makes it inherently lightweight and easy to integrate into existing single-frame models.
((a))
((b))
Figure 2: (a) The procedure of basis decomposition, where $\vec{b_{1}}$, $\vec{b_{2}}$, and $\vec{b_{3}}$ are the basis vectors. In our study, we regard the results obtained by difference kernels or frame-difference methods as our basis vectors. (b) The illustration of the proposed module for signal recovery and task-relevant feature enhancement. First, we collect both task-relevant and task-irrelevant basis features into a basis pool according to the specific tasks. Then, the input is processed in two parallel branches for recovery and enhancement. The upper branch indicates the recovery process, where the input is reconstructed through basis decomposition. The lower branch outlines the enhancement procedure, which selects the task-relevant basis features to reconstruct the input.
II-C Difference-based Methods
In recent years, there have been increasingly difference-based methods in various methods [Su2025Rapid] to strengthen the detailed information in spatial and temporal domains.
In the spatial domain, researchers often use 2D convolution to extract spatial information. However, standard convolution is not sufficiently effective in obtaining the detailed information crucial for IRSTD. Therefore, there are some methods based on the classical descriptors, like Local Binary Pattern (LBP) [Ahonen2006LBP] and Gabor filters[JAIN1991Gaborfilter], to strengthen details [Xu2017LBPConv, luan2018GraborConv]. Building on this, central difference convolutions (CDCs) are proposed [yu2020CDC]. Chen et al. [Chen2024DEANet] parallel the vanilla convolutions and difference convolutions to obtain various features. Kim et al. [Kim2021Deformable] propose the Deformable Kernel Network, with adaptive neighbors and their corresponding weights.
In the temporal domain, some works choose to use the difference results between frames to obtain motion information[Yan2023STDMANet]. Besides, Geng et al. [GENG20253DCDC] propose the 3D central difference separable convolution and introduce the motion details into the difference convolution.
All of the studies demonstrated the strong power on highlighting the details and obtaining the motion information. Therefore, our method is justified.
@@ III-A Preliminaries
III-A1 The procedure of basis decomposition
@@
As shown in the upper part of Figure 2(b), a vector $\vec{x}\in\mathbb{R}^{c_{\text{ful}}}$ can be recovered by $c_{ful}$ orthonormal vectors:
@@
where $s_{i}$ denotes the coefficient of $\vec{x}$ projected onto the basis vector $\vec{b}_{i}$. $\top$} is the transposition operation, and $c_{\text{ful}}$ is the dimension of the input vector. Additionally, dimdim denotes the dimensionality of the vector $\vec{x}$.
((a))
((b))
Figure 3: Robustness of our methods. (a) Illustration of robustness to data value variations. The representation is constructed based on normalized relative differences between neighboring elements, making it insensitive to uniform shifts or scale changes in the input values. (b) Illustration of rotation invariance. Local patterns under different input directions are aligned to an orientation, resulting in identical weight arrangements for rotated versions of the same pattern.
Furthermore, the result can be constructed solely from the selected basis elements and the irrelevant basis features are removed, thereby enhancing the task-relevant features, as illustrated in the lower portion of Figure 2(b). In this way, we can strengthen the important features by reducing interference from non-essential information.
@@
III-A2 Relaxing Constraints:
Basis vectors must be both normalized and orthogonal. As shown in Table I, our method does not require the basis features to be strictly orthogonal, because forcing $\{\vec{b}_{k}\}$ to be orthogonal would require an additional decorrelation step, which would mix these simple local difference patterns into more global filters and destroy their physical interpretation. By contrast, normalization plays essential roles, which removes the scale ambiguity between a basis vector and its coefficient, prevents basis features with larger norms from dominating the decomposition, and improves robustness. Thus, the orthogonality constraints are relaxed, while the normalization requirement is retained.
Figure 4: Overview of the BDM.
III-A3 Difference Decomposition
When it comes to IRSTD, where detection relies more on local contrast than on absolute intensity, it is natural to construct features from difference information. Consequently, we treat the difference features, obtained by difference kernels or frame difference methods, as basis feature vectors and combine them to form $\vec{y}$ and approximate $\vec{x}$. Ultimately, $\vec{y}$, the fusion of difference features, enhances the high-frequency components that contain information about infrared small targets. In our module, the decomposition coefficients are not learned as fixed parameters. Instead, they are computed dynamically from the current input. Thus, the coefficients adapt to each sample and each location, rather than being shared globally across all inputs. Difference decomposition has two main properties:
  • 1.
    We begin by performing difference extraction and normalization to obtain the basis vectors from the relative differences, as illustrated in Figure 3(a). This approach to difference decomposition mitigates the impact of overall intensity changes in features, as each basis is defined based on differences rather than absolute values.
  • 2.
    The output is formed by combining the basis features with their corresponding coefficients, where the coefficients are computed from the input features and the basis feature group. Consequently, if both the input features and the basis feature group remain unchanged, the output of our module will also remain unchanged. For example, under spatial difference decomposition, even when the input undergoes a rotation when the basis features remain fixed, the output retains the origin result, as shown in Figure 3(b).
Therefore, due to the difference decomposition method, it offers more robust and effective performance compared to conventional techniques, such as CDCs and frame-difference methods.